hartnn (hartnn):

Fermat's Little Theorem. a^{p-1}=(corresponds to)1mod p using this, we have remainder when 3^100 is divided by 101 as 1. Is this correct ? Can i have few more examples where this theorem is applied but not so directly... @mukushla

hartnn (hartnn):

@siddhantsharan , @sauravshakya

hartnn (hartnn):

yes, that was direct use of the theorem....

OpenStudy (anonymous):

i think we just use it directly

hartnn (hartnn):

any other interesting topic or theorem in modular arithmetic ?

OpenStudy (anonymous):

Wilson's theorem is interesting

hartnn (hartnn):

(n-1)!=-1(mod n) how do we use it ?

OpenStudy (anonymous):

If \(p\) is a prime, then \((p-1)!+1\) is a multiple of \(p\), that is\[(p-1)! \equiv -1 \ \ \text{mod} \ p\]

hartnn (hartnn):

i will be glad if someone posts a link for a good reference in this topic...

OpenStudy (anonymous):

FLT is also a^p = a (mod p) Show that the two formulations are equivalent.

hartnn (hartnn):

divide by a on both sides? maybe.....

OpenStudy (anonymous):

It's an iff type, so both ways......

hartnn (hartnn):

idk....maybe multiply both sides by a to get a^p = a (mod p) but is it correct to do so ?

OpenStudy (anonymous):

On the right track, sort of...

OpenStudy (anonymous):

a^p = a mod p, then if a not = 0 mod p then can cancel by a to get a^(p-1) = 1 mod p That's the first part...

hartnn (hartnn):

same for other part, right ?

OpenStudy (anonymous):

Yes, more or less.....

OpenStudy (anonymous):

FLT is also used in some combinatorial problems....

hartnn (hartnn):

u have some practice/examples with those ?

OpenStudy (anonymous):

OK, you can try this one: Say you have enough beads in n colours, how many different necklaces consisting of p beads can be made, where p is prime?

hartnn (hartnn):

(mod(n/p))!

OpenStudy (anonymous):

I guess I should tell you that you will need to think a bit in order to work it out, it is not straightforward....

OpenStudy (anonymous):

Or I can tell you the answer and you can try to work it out.....

hartnn (hartnn):

let me try....after few minutes, u give answer...

OpenStudy (anonymous):

I think u need more than a few minutes....:-)

hartnn (hartnn):

ok, whats the answer ?

OpenStudy (anonymous):

(n^p-n)/2p + (n^((p+1)/2) -n)/2 + n

hartnn (hartnn):

OMG!

OpenStudy (anonymous):

It's an integer so the first term still incorporates FLT...

OpenStudy (anonymous):

Do you want links only about FLT/Wilsons or congruences/modular stuff as well....

hartnn (hartnn):

modular stuff as well...

OpenStudy (anonymous):

Similar this http://www.math.sc.edu/~filaseta/gradcourses/Math780notes.pdf Or more detailed....

hartnn (hartnn):

thanks, i'll go through it......and will ask u for any doubts.